OpenStudy (anonymous):

A partial solution to the equation 2(3x + 6) = 24 is given below. Choose the step that is missing from the solution. 2(3x + 6) = 24 ____________ 6x = 12 x = 2

OpenStudy (anonymous):

2(3x+6)=24 6x+12=24 <---- 6x=24-12=12 x=2

OpenStudy (anonymous):

thank you that's what i get but i was not sure

OpenStudy (anonymous):

your welcome